# A call for frugal modelling: two case studies involving molecular spin dynamics

### 24 Jan 2024 v1

@@ Conclusion
We presented here two particular examples illustrating an extremely common situation: one can often choose either to solve a different model or to solve it via a different implementation, and obtain significant savings in carbon footprint without making any significant sacrifices in scientific yield. Currently this is mostly overlooked, and if anything the most expensive methods tend to enjoy a higher prestige and are considered more trustworthy. Often, this means that wasteful methods allow for easier or better publishing venues. In our case, the research presented in Liu et al 31 was originally submitted with the frugal method, but during the refereeing process we switched to the wasteful method as a requisite to acceptance. As a community we need to do better, and the main factor is choosing affordable models to solve problems, with a minor but sizeable contribution of finding an inexpensive way to implement these models. Crucially, we need to beware of the Jevons’ paradox or rebound effect, meaning a more efficient method generally leads to an increased usage which often overshoots the savings. Thus, we call for frugality rather than for efficiency. Just as we consider the ethical repercussions of animal experimentation, or dealing with patient data, eventually we will need to include the risk of carbon footprint wastefulness as an ethical concern in research.
@@
The data reported in this work and in the Supplementary Information are available from the corresponding author upon reasonable request.
@@ Acknowledgements
A.G.A. has been supported by the COST Action MolSpin on Molecular Spintronics (Project 15128), by the European Union (EU) Programme Horizon 2020 (FATMOLS project), and by the Generalitat Valenciana (GVA) CIDEGENT/2021/018 grant. A.G.A. thanks grant PID2020-117177GB-I00 funded by MCIN/AEI/10.13039/501100011033 (both MCIN/AEI grants co-financed by FEDER funds). This study is part of the Quantum Communication programme and was supported by grant PRTR-C17.I1 funded by MCIN/AEI/10.13039/501100011033 and European Union NextGenerationEU/PRTR, and by GVA (QMol COMCUANTICA/010).

### 20 Feb 2024 v2

@@ Conclusion
We presented here two particular examples illustrating an extremely common situation: when confronted to a calculation with a large carbon footprint, one can often choose either to solve a different model or to solve the same model via a different implementation, and obtain significant savings in carbon footprint without making any significant sacrifices in scientific yield. Currently this is mostly overlooked, and if anything the most expensive methods tend to enjoy a higher prestige and are considered more trustworthy. Often, this means that wasteful methods allow for easier or better publishing venues. In our case, the research presented in Liu et al 35 was originally submitted with the frugal method, but during the refereeing process we were requested to switch to the method that is in principle more exact but which we show here is actually wasteful in this case. As a community we need to do better, and the main factor is choosing affordable models to solve problems, with a minor but sizeable contribution of finding an inexpensive way to implement these models. Crucially, we need to beware of the Jevons’ paradox or rebound effect, 37 meaning a more efficient method, if not coupled to resource consciousness, by itself leads to an increased usage which often overshoots the savings. Thus, we call herein for ”frugality” rather than for ”efficiency”. More generally, just as we consider the ethical repercussions of animal experimentation, or dealing with patient data, eventually we will need to include the risk of carbon footprint wastefulness as an ethical concern in research.
@@
The data reported in the first part of this work are available as part of the Supplementary Information.For the second part, all custom data generated and employed for this study are available at https://github.com/gerlizg/STOSS-MATLAB.
CODE AVAILABILITY
For the second section of this paper, the original code (python version) named STOSS (for STOchastic Spin Simulator) is available at https://github.com/gerlizg/STOSS. The new version of the program using Matlab and the instructions to reproduce all the results could be found in the Supporting Information Section S2 and at https://github.com/gerlizg/STOSS-MATLAB.
@@ Acknowledgements
A.G.A. has been supported by the European Union (EU) Programme Horizon 2020 (FATMOLS project), and by the Generalitat Valenciana (GVA) CIDEGENT/2021/018 grant. A.G.A. thanks grant PID2020-117177GB-I00 funded by MCIN/AEI/10.13039/501100011033 (co-financed by FEDER funds). This study is part of the Quantum Communication programme and was supported by grant PRTR-C17.I1 funded by MCIN/AEI/10.13039/501100011033 and European Union NextGenerationEU/PRTR, and by GVA (QMol COMCUANTICA/010).